OpenStudy (anonymous):

WRITE THE ANSWER IN slope-intercept and and standard form! parallel to y=2/3x - 5 and through (2,7) HELP!!!!

OpenStudy (anonymous):

Same slope, different y intercept

OpenStudy (anonymous):

SO SLOPE IS 2/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and to find the y intercept would i use the points (2,7) to get 2/7?

OpenStudy (anonymous):

or would i have to plug variables into an equation?

OpenStudy (anonymous):

you wouldn't use the point (2, 7) that's the point your trying to hit

OpenStudy (anonymous):

so how would i solve for the y intercept?

OpenStudy (anonymous):

my equation so far is: y=1/3x + b?

OpenStudy (anonymous):

I think that is 5.67

OpenStudy (anonymous):

y-7=2/3(x-2) y=2/3x+17/3

OpenStudy (anonymous):

what would that be?

OpenStudy (anonymous):

standard form?

OpenStudy (anonymous):

sorry slope intercept form, right?

OpenStudy (anonymous):

yep slope intercept form y=mx+b

OpenStudy (anonymous):

standard is 2x+3y=-17!!!